OpenStudy (anonymous):

what is the slope of a line that is perpendicular to the line represented by the equation 3x+4y=12 ???

jimthompson5910 (jim_thompson5910):

your first task is to isolate y

jimthompson5910 (jim_thompson5910):

3x+4y=12 3x+4y-3x = 12 - 3x 4y = -3x + 12 y = -3x/4 + 12/4 y = (-3/4)x + 3 This last equation is in the form y = mx+b where m = -3/4 b = 3

jimthompson5910 (jim_thompson5910):

So the slope of y = (-3/4)x + 3 and 3x+4y=12 is m = -3/4

jimthompson5910 (jim_thompson5910):

to find the slope of any perpendicular line to this one, you simply flip the fraction and flip the sign flip the fraction: -3/4 -----> -4/3 flip the sign: -4/3 -----> +4/3 or just 4/3 So the slope of any perpendicular line to 3x+4y=12 is 4/3

OpenStudy (anonymous):

slope of the line ax+by=c is given by -(coeifficient of x/coefficient of y)

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw